OpenStudy (anonymous):

Physics please help A plane travels 350 km south along a straight path with an average velocity of 125 km/h to the south. The plane has a 30 minute layover, and then travels 220 km south with an average velocity of 115 km/h to the south. What is the average velocity for the trip? a.120 km/hr to the south b.110 km/hr to the south c. 121 km/hr to the south d. need more information

OpenStudy (amistre64):

distance traveled, divided by the amount of time it took to get there.

OpenStudy (theeric):

And you'll find the time of the flights using the same calculation on each part of the trip, and don't forget the 30 minute layover part! Remember the definition of velocity, and put that into math! That'll be the formula you can use to find the times. Then you add up the distances to get the total distance. Then you add up the times to get the total time. Then you use the equation from earlier, pretty much. Good luck!

OpenStudy (anonymous):

Since average speeds of the trip are already given, just simply add the given average speeds then divide it by two. So the answer is 120 km/h south. There is no given time, then the formula distance divided by time cannot be used.

OpenStudy (anonymous):

ans is 121.27 kmph ~ 121kmph--- s1=350 km + s2=220km =570 km t1=350/125 = 2.8 hr, t2=220/115 hr= 1.9 hr. V[avg]= 350+220/4.7 hr= 121.27 kmph

OpenStudy (amistre64):

\[\frac{350}{120}~\frac{km}{km/hr}=\frac{35}{12}hr~:~\text{so a time is implicitly given for each leg of the trip}\]

OpenStudy (amistre64):

as long as you use the correct information given and not a typo :)